The equilibrium constant Kc for a reaction at 842℃ is 7.90×10^-3 . Find Kp for this reaction at same temperature.
  1/2 F2(g) F(g)

Grade:11

1 Answers

Sangeeta
5 years ago
We know Kp=Kc×(RT)∆ng.  ------1
   , where R=0.0821lit atm. 
               T=1115K
Given that Kp=7.9×10--3
∆ng=1-1/2=1/2
On puting values in 1
We get Kp=7.56 × 10--2